The INORDER traversal output of a binary tree is U,N,I,V,E,R,S,I,T,Y and the POSTORDER traversal output of the same tree is N,U,V,R,E,T,I,S,I,Y

 a. The INORDER traversal output of a binary tree is U,N,I,V,E,R,S,I,T,Y and the POSTORDER traversal output of the same tree is N,U,V,R,E,T,I,S,I,Y. Construct the tree and determine the output of the PREORDER traversal output.  

 b. One main difference between a binary search tree (BST) and an AVL (Adelson-Velski and Landis) tree is that an AVL tree has a balance condition, that is, for every node

 in the AVL tree, the height of the left and right subtrees differ by at most 1. Starting with an empty BST and AVL tree, insert elements into the two trees with the following keys:

 27, 32, 89, 45, 10, 63, 57, 17

 c. Based on what you have done for Question 1b, what is the big-O (worst case) complexity of the total time required to build a binary search tree (BST) consisting of n nodes? Explain your answer. Answer without explanation gains no mark. (Hint.

 A tricky question. The question asks for worst case complexity. Think properly.)

 d. Similarly, from what you have done for Question 1b, what is the big-o (worst case) complexity of the total time required to build an AVL tree consisting of n nodes? Explain your answer. Answer without explanation gains no mark. (Hint. The question asks for worst case complexity.)
Answer #1

Solution:

(a) & (b):

Explanation:

(c)

Explanation:

=>Worst case time complexity of building BST(binary search tree) = O(n^2) where n is the number of elements to be inserted into the empty binary search tree.

Calculations:

=>Searching for a key takes time in worst case = O(n) as height of tree can be O(n)

=>After searching the key insertion takes time = O(1)

=>Total time for one key in worst case = O(n) + O(1)

=>Total time for one key in worst case = O(n)

=>Total time for n keys = n*O(n)

=>Total time for n keys = O(n^2)

(d)

Explanation:

=>Worst case time complexity of AVL tree insertion for n elements = O(nlogn)

Calculations:

=>Searching for a key takes time in worst case = O(logn) as height of tree can be O(logn)

=>After searching the key insertion takes time = O(1)

=>Total time for one key in worst case = O(logn) + O(1)

=>Total time for one key in worst case = O(logn)

=>Total time for n keys = n*O(logn)

=>Total time for n keys = O(nlogn)